### v2

@@ 3.2.1 Intrinsic-CFA
Intrinsic-CFA enhances the feature of query triplet by replacing its visual features of entities (i.e., subject and object). It consists of two steps: entity category selection and entity feature selection and replacement (cf.. Fig. 5(a)). During training, we randomly select a query triplet from a batch of images, and randomly select one of its entity features as input. Then we put it into the Intrinsic-CFA module (cf.. Fig. 4). Next, we detailedly introduce each step.
Entity Category Selection. This step is used for determining the category to which the entity feature of the query triplet is replaced. Firstly, we propose a novel relation-conditioned entity clustering that organizes entity categories based on their semantic role in a triplet. The framework consists of the following procedures.
(1) Predicate Clustering. To maintain consistency in entity substitutions, we first group predicates into several clusters. The clustering process utilizes three similarity measures:
  • Pattern Similarity: This metric quantifies the degree of functional overlap between predicates by analyzing the frequency with which they govern similar subject-object pairs. Formally, for two predicate categories $c_{p_{i}}$ and $c_{p_{j}}$, pattern similarity is defined as:
    $\small Sim_{p}(c_{p_{i}},c_{p_{j}})=\frac{|L_{so}|}{{d(c_{p_{i}})+d(c_{p_{j}})-|L_{so}|}},$ (6)
    where $|L_{so}|$ denotes the number of common subject-object associations and $d(c_{p_{i}})$ denotes the sample number with the predicate category $c_{p_{i}}$ in the dataset.
  • Context Similarity: This measure evaluates the extent to which two predicates appear in similar environmental contexts by comparing their co-occurring entities:
    $\small Sim_{c}(c_{p_{i}},c_{p_{j}})=\frac{|L_{e}|}{{d_{co}(c_{p_{i}})+d_{co}(c_{p_{j}})-|L_{e}|}},\$ (7)
    where $|L_{e}|$ denotes the number of entity categories that co-occur with both $c_{p_{i}}$ and $c_{p_{j}}$ in the same image. The term $d_{co}(c_{p_{i}})$ represents the number of images in which predicate category $c_{p_{i}}$ appears.
  • Semantic Similarity: To capture semantic alignment, we compute the Euclidean distance between two predicate categories in a pre-trained embedding space:
    $Sim_{s}(c_{p_{i}},c_{p_{j}})=||\bm{e}_{p_{i}}-\bm{e}_{p_{j}}||_{2}.$ (8)
The final similarity score for predicate clustering is a weighted sum of these three measures.
(2) Predicate Category Merging. Once predicate clusters are established, associated subject and object categories within each cluster are merged to ensure consistency. Overlapping subject categories among clustered predicates are unified, guaranteeing consistent treatment of interchangeable subjects across relational structures. Likewise, object categories governed by these clustered predicates are merged, preserving the relational and contextual properties of entity substitutions.
For instance, if the predicates standing on and sitting on belong to the same cluster, their subject categories (e.g., dog, cat, person) and object categories (e.g., chair, floor, bed) are merged accordingly.
(3) Entity Category Clustering. After merging, we further cluster entity categories within the same predicate clusters to identify interchangeable entities. The clustering process follows a methodology similar to predicate clustering and employs three similarity measures. Fig. 6 shows an example of pattern and context similarity between cat and dog:
  • Pattern Similarity: This metric evaluates the extent to which two entity categories participate in similar relational structures. It is computed as:
    $Sim_{p}(c_{e_{i}},c_{e_{j}})$ $=\frac{|L_{sp}|}{{d_{out}(c_{e_{i}})+d_{out}(c_{e_{j}})-|L_{sp}|}}$ (9)
    $+\frac{|L_{po}|}{{d_{in}(c_{e_{i}})+d_{in}(c_{e_{j}})-|L_{po}|}},$
    where $|L_{sp}|$ and $|L_{po}|$ represent the number of common subject-predicate and predicate-object associations. Terms $d_{in}(c_{e_{i}})$ and $d_{out}(c_{e_{i}})$ denote the number of incoming and outgoing edges of entity category $c_{e_{i}}$.
  • Context Similarity: Similarly, this similarity measures the context’s overlap of two entity categories in images:
    $\small Sim_{c}(c_{e_{i}},c_{e_{j}})=\frac{|L_{e}|}{{d_{co}(c_{e_{i}})+d_{co}(c_{e_{j}})-|L_{e}|}}.\$ (10)
  • Semantic Similarity: As with predicates, it’s the distance between two entity semantic embeddings:
    $Sim_{s}(c_{e_{i}},c_{e_{j}})=||\bm{e}_{e_{i}}-\bm{e}_{e_{j}}||_{2}.$ (11)
Entire clustering algorithm and results are in the appendix.
@@
Entity Feature Selection and Replacement. This step selects an entity feature from the feature bank based on the selected entity category and replaces it with the query triplet. Specifically, we first update a new entity category of query triplet (e.g., cat-laying on-bed). Then, we select all triplet features with the same category as query triplet from feature bank as candidates (cf., Fig. 5(a)). And we utilize spatial restriction to filter unreasonable triplet features.
Spatial Restriction. In real augmentation, even if the two entity categories appear to be interchangeable, the positions of the subject and object after replacement may not make sense. For example, in VG dataset (cf.. appendix), though category hat and shoe are in the same cluster, directly replacing hat with shoe would place the shoe on the head and make it contrary to common sense.
To further ensure the rationality of the replacement of the two entity features, we use the cosine similarity of the relative spatial position of subject-object of the query triplet and the selected triplet from the feature bank as restriction:
@@
Highlights. In the original intrisic-CFA of CFA Li et al. (2023a), we clustered entities uniformly, ignoring predicate-driven semantic discrepancy (e.g., kite and umbrella are replaceable for holding, but an umbrella is ill-suited for flying in the sky where a kite fits). In this paper, we devise a relation-conditioned entity clustering, allowing more accurate, predicate-aware entity groupings. This approach, for instance, discerns the similarity of kite and umbrella for holding versus their distinct roles for flying in, yielding more precise entity selection.
@@ 3.2.2 Extrinsic-CFA
Extrinsic-CFA aims to enrich the feature of the query triplet through the extrinsic information (i.e., the context formed by all the entities of an image) of context triplets. Different from the intrinsic-CFA, we select context triplet during training for computation efficiency. Extrinsic-CFA includes three steps: context triplet selection, query triplet feature selection, and adaptive feature augmentation (cf., Fig. 5(b)). Then, we detail each step of extrinsic-CFA.
@@
Query Triplet Feature Selection. This step selects a query triplet with its features for the feature augmentation. Specifically, we select triplets with the same subject-object categories (e.g., pillow-bed) as the context triplet (e.g., pillow-on-bed) from feature bank as the candidate query triplets. Then, we implement the same spatial restriction as intrinsic-CFA to filter unreasonable triplets. Finally, we randomly select one of the reasonable triplets as the query triplet (e.g., pillow-laying on-bed), and its features are utilized for augmentation (cf.. Fig. 5(b)).
Adaptive Feature Augmentation. To enhance the extrinsic features of tail predicate triplets while concurrently mitigating the impact on other triplets of the original image, we devise a novel adaptive mixup operation. This operation blends a selected query triplet (e.g., pillow-laying on-bed) with a context triplet (e.g., pillow-on-bed), as depicted in Fig. 5(b). The adaptive mixup is formulated as follows:
[…]
$\tilde{\bm{v}}_{s}=\theta\bm{v}_{s}+(1-\theta){{\bm{v}}_{s}^{\prime}},$
$\tilde{\bm{v}}_{o}=\theta{\bm{v}}_{o}+(1-\theta){{\bm{v}}_{o}^{\prime}},$
$\tilde{\bm{u}}=\theta\bm{u}+(1-\theta){{\bm{u}}^{\prime}},$
where $\bm{e}_{p}$ and $\bm{e}_{p}^{\prime}$ denote the semantic embedding of relation category in context triplet (e.g., on) and selected query triplet (e.g., laying on), respectively. The parameters $\alpha$ is used to control the value range of $\theta$. $\bm{v}_{s}^{\prime}$ ($\bm{v}_{s}$) and $\bm{v}_{o}^{\prime}$ ($\bm{v}_{o}$) denote the visual features of the subject and object of the query triplet (context triplet), respectively. The $\bm{u}^{\prime}$ ($\bm{u}$) denotes the visual feature of the union box of the query triplet (context triplet). Similarly, the predicate target ground-truth $r$ of the context triplet is mixed:
@@
After the mixup operation, the query triplet feature can be enhanced in context modeling (cf.. Fig. 4) by leveraging the extrinsic features of the context triplet.
Highlights. In CFA Li et al. (2023a), we employed a fixed $\theta$ for feature augmentation. It fails to adapt to varying semantic similarities between query and context triplets, risking under-enhancement of tail predicate triplets or feature corruption via suboptimal contextual blending. In this paper, we propose an adaptive feature augmentation strategy, enabling the mixup coefficient $\theta$ to dynamically adjust based on predicate similarity between query and context triplets. This facilitates targeted enrichment of tail predicates using relevant contexts while preventing feature distortion from less aligned ones, ultimately boosting model robustness.
3.3 Image-level CFA
Although the above mentioned methods effectively augment the diversity of both extrinsic and intrinsic features of predicates, the constructed feature space is still limited in the dataset domain. In order to maximally infer the real feature space of predicates and learn the true data distribution, we also propose an image-level augmentation method that leverages large generative models Li et al. (2023b) to create new visual appearance of relation triplets. Image-level CFA includes three steps: context triplet and query triplet selection, augmented triplet image generation, and quality estimation (cf., Fig. 5(c)). Then, we detail each step of Image-level CFA.
Context Triplet and Query Triplet Selection. For each image, this step selects context triplets and query triplet to be qugmented for image-level augmentation. We select context triplets with foreground predicate categories to provide extrinsic information. We apply the same sampling strategy as the way of extrinsic-CFA in Sec. 3.2.2, i.e., we randomly sample the context triplets by using probability $p=(\eta-{\eta_{r}})/\eta\times\gamma$ during training, where $\gamma$ is hyperparameter. We also apply the same strategy as the way of extrinsic-CFA in Sec. 3.2.2 to select query triplet, but we only use its categories and grounding information of entities and predicate. Then we send them with context triplets together to generate new visual images.
Augmented Triplet Image Generation. This step is to generate new image with context and query triplets. The main challenge is to ensure the rationality of generated image with query triplet. Taking the power of recent generative models, e.g., GLIGEN Li et al. (2023b), we generate new samples conditioned on more input information apart from text to control the generation process, e.g., bounding boxes and reference images.
Specifically, for each image $I$, suppose that the subject and object of the selected query triplet is $i$ and $j$, and the corresponding bounding boxes are $\bm{b}_{i},\bm{b}_{j}\in\mathcal{B}$ respectively. Our aim is to generate images with new visual appearance of entity $i$ and $j$ conditioned on the context information. We send the instruction $\bm{g}$ to the grounded text-to-image model GLIGEN Li et al. (2023b), which is a composition of reference image and query triplet information:
$\bm{g}=(\bm{e}_{i},\bm{e}_{j},r_{ij},I),\text{with}\;\bm{e}_{i}=(\bm{b}_{j},o_{i}\in\mathcal{O})$ (15)
where $\bm{e}_{i}$ are the information of entity $i$ with its category and bounding box, and $r_{ij}\in\mathcal{R}$ is the triplet predicate category. After pretrained with multiple input modalities besides text input, GLIGEN Li et al. (2023b) is able to generate improved quality image with grounding conditions. Some generated samples are displayed in Fig. 8.
Quality Estimation. After generating the new visual appearance of entities, we have to check their quality and make sure their category are remained consistent. We send the cropped regions of generated entities and triplets into the pretrained CLIP model Radford et al. (2021) to predict the classification scores. Only the correctly classified samples will be selected as the augmented datas to participate in the training process of SGG models.
@@ 3.4 Training Objectives
Contrastive Loss. Due to the pattern changes between the entity features after mixup and original entity features, using only XE losses may result in performance drops in entity classification. To maintain the discriminative entity features after mixup, we further apply a contrastive loss Chen et al. (2020):
@@
Motifs  Zellers et al. (2018)${}_{\textit{CVPR'18}}$ 16.5 17.8 65.5 67.2 41.8 8.7 9.3 39.0 39.7 24.2 5.5 6.8 32.1 36.9 20.3
VCTree Tang et al. (2019)${}_{\textit{CVPR'19}}$ 17.1 18.4 65.9 67.5 42.2 10.8 11.5 45.6 46.5 28.6 7.2 8.4 32.0 36.2 20.9
Transformer Vaswani et al. (2017)${}_{\textit{NIPS'17}}$ 17.9 19.6 63.6 65.7 41.7 9.9 10.5 38.1 39.2 24.4 7.4 8.8 30.0 34.3 20.1
BGNN Li et al. (2021b)${}_{\textit{CVPR'21}}$ 30.4 32.9 59.2 61.3 45.9 14.3 16.5 37.4 38.5 26.7 10.7 12.6 31.0 35.8 22.5
Motifs+PCPL Yan et al. (2020)${}_{\textit{ACMMM'20}}$ 24.3 26.1 54.7 56.5 40.4 12.0 12.7 35.3 36.1 24.0 10.7 12.6 27.8 31.7 20.7
Motifs+DLFE Chiou et al. (2021)${}_{\textit{ACMMM'21}}$ 26.9 28.8 52.5 54.2 40.6 15.2 15.9 32.3 33.1 24.1 11.7 13.8 25.4 29.4 20.1
Motifs+BPL-SA Guo et al. (2021)${}_{\textit{ICCV'21}}$ 29.7 31.7 50.7 52.5 41.2 16.5 17.5 30.1 31.0 23.8 13.5 15.6 23.0 26.9 19.8
Motifs+NICE Li et al. (2022a)${}_{\textit{CVPR'22}}$ 29.9 32.3 55.1 57.2 43.6 16.6 17.9 33.1 34.0 25.4 12.2 14.4 27.8 31.8 21.6
Motifs+IETrans Zhang et al. (2022)${}_{\textit{ECCV'22}}$ 30.9 33.6 54.7 56.7 44.0 16.8 17.9 32.5 33.4 25.2 12.4 14.9 26.4 30.6 21.1
Motifs+CFA Li et al. (2023a)${}_{\textit{ICCV'23}}$ 35.7 38.2 54.1 56.6 46.2 17.0 18.4 34.9 36.1 26.6 13.2 15.5 27.4 31.8 22.0
Motifs+DPL Jeon et al. (2024)${}_{\textit{ECCV'24}}$ 33.7 37.4 54.4 56.3 45.5 18.5 20.1 32.6 33.8 26.3 13.0 15.6 24.5 28.7 20.5
Motifs+MCFA (ours) 35.8 38.3 54.3 56.8 46.3 17.9 19.4 35.5 36.7 27.4 14.0 16.3 27.1 31.6 22.3
VCTree+PCPL Yan et al. (2020)${}_{\textit{ACMMM'20}}$ 22.8 24.5 56.9 58.7 40.7 15.2 16.1 40.6 41.7 28.4 10.8 12.6 26.6 30.3 20.1
VCTree+DLFE Chiou et al. (2021)${}_{\textit{ACMMM'21}}$ 25.3 27.1 51.8 53.5 39.4 18.9 20.0 33.5 34.6 26.8 11.8 13.8 22.7 26.3 18.7
VCTree+BPL-SA Guo et al. (2021)${}_{\textit{ICCV'21}}$ 30.6 32.6 50.0 51.8 41.3 20.1 21.2 34.0 35.0 27.6 13.5 15.7 21.7 25.5 19.1
VCTree+NICE Li et al. (2022a)${}_{\textit{CVPR'22}}$ 30.7 33.0 55.0 56.9 43.9 19.9 21.3 37.8 39.0 29.5 11.9 14.1 27.0 30.8 21.0
VCTree+IETrans Zhang et al. (2022)${}_{\textit{ECCV'22}}$ 30.3 33.9 53.0 55.0 43.1 16.5 18.1 32.9 33.8 25.3 11.5 14.0 25.4 29.3 20.1
VCTree+CFA Li et al. (2023a)${}_{\textit{ICCV'23}}$ 34.5 37.2 54.7 57.5 46.0 19.1 20.8 42.4 43.5 31.5 13.1 15.5 27.1 31.2 21.7
VCTree+DPL Jeon et al. (2024)${}_{\textit{ECCV'24}}$ 35.3 37.9 54.0 55.8 45.8 19.2 21.3 32.6 33.9 26.8 13.2 15.9 27.0 31.4 21.9
VCTree+MCFA (ours) 35.4 38.2 54.0 56.9 46.1 21.0 23.3 41.8 42.9 32.3 13.4 15.8 27.2 31.6 22.0
Transformer+IETrans Zhang et al. (2022)${}_{\textit{ECCV'22}}$ 30.8 34.5 51.8 53.8 42.7 17.4 19.1 32.6 33.5 25.7 12.5 15.0 25.5 29.6 20.7
Transformer+CFA Li et al. (2023a)${}_{\textit{ICCV'23}}$ 30.1 33.7 59.2 61.5 46.1 15.7 17.2 36.3 37.3 26.6 12.3 14.6 27.7 32.1 21.7
Transformer+MCFA (ours) 32.4 34.4 58.9 61.0 46.7 16.4 17.6 36.7 37.9 27.2 12.6 15.1 27.4 32.0 21.8
Table 1: Performance (%) of the SOTA trade-off SGG models on VG Krishna et al. (2017). “Mean” is the average of mR@50/100 and R@50/100.
@@
Motifs+TDE Tang et al. (2020)${}_{\textit{CVPR'20}}$ 18.5 25.5 29.1 9.8 13.1 14.9 5.8 8.2 9.8
Motifs+CogTree Yu et al. (2021)${}_{\textit{IJCAI'21}}$ 20.9 26.4 29.0 12.1 14.9 16.1 7.9 10.4 11.8
Motifs+RTPB Chen et al. (2022)${}_{\textit{AAAI'22}}$ 28.8 35.3 37.7 16.3 20.0 21.0 9.7 13.1 15.5
Motifs+PPDL Li et al. (2022c)${}_{\textit{CVPR'22}}$ 27.9 32.2 33.3 15.8 17.5 18.2 9.2 11.4 13.5
Motifs+GCL Dong et al. (2022)${}_{\textit{CVPR'22}}$ 30.5 36.1 38.2 18.0 20.8 21.8 12.9 16.8 19.3
Motif+HML Deng et al. (2022)${}_{\textit{ECCV'22}}$ 30.1 36.3 38.7 17.1 20.8 22.1 10.8 14.6 17.3
Motifs+CFA Li et al. (2023a)${}_{\textit{ICCV'23}}$ 31.5 39.9 43.0 17.3 20.9 22.4 11.2 15.3 18.1
Motifs+MCFA (ours) 31.8 40.3 43.5 17.7 21.3 22.7 11.9 15.8 18.6
VCTree+TDE Tang et al. (2020)${}_{\textit{CVPR'20}}$ 18.4 25.4 28.7 8.9 12.2 14.0 6.9 9.3 11.1
VCTree+CogTree Yu et al. (2021)${}_{\textit{IJCAI'21}}$ 22.0 27.6 29.7 15.4 18.8 19.9 7.8 10.4 12.1
VCTree+RTPB Chen et al. (2022)${}_{\textit{AAAI'22}}$ 27.3 33.4 35.6 20.6 24.5 25.8 9.6 12.8 15.1
VCTree+PPDL Li et al. (2022c)${}_{\textit{CVPR'22}}$ 29.7 33.3 33.8 20.3 21.8 22.4 9.1 11.3 13.3
VCTree+GCL Dong et al. (2022)${}_{\textit{CVPR'22}}$ 31.4 37.1 39.1 19.5 22.5 23.5 11.9 15.2 17.5
VCTree+HML Deng et al. (2022)${}_{\textit{ECCV'22}}$ 31.0 36.9 39.2 20.5 25.0 26.8 10.1 13.7 16.3
VCTree+CFA Li et al. (2023a)${}_{\textit{ICCV'23}}$ 31.6 39.2 42.5 21.5 26.3 28.3 10.8 15.1 17.9
VCTree+MCFA (ours) 32.3 40.2 43.2 21.7 27.2 29.2 11.0 15.3 18.0
Transformer+CogTree Yu et al. (2021)${}_{\textit{IJCAI'21}}$ 22.9 28.4 31.0 13.0 15.7 16.7 7.9 11.1 12.7
Transformer+HML Deng et al. (2022)${}_{\textit{ECCV'22}}$ 27.4 33.3 35.9 15.7 19.1 20.4 11.4 15.0 17.7
Transformer+CFA Li et al. (2023a)${}_{\textit{ICCV'23}}$ 31.2 38.6 41.5 17.2 20.9 22.7 10.6 15.0 17.9
Transformer+MCFA (ours) 31.4 38.9 42.2 17.5 21.5 23.1 10.8 15.1 18.1
@@
+MCFA 28.6 / 30.4 16.6 / 17.2 13.5 / 15.6
Table 3: Performance (%) of the SGG models on GQA Hudson and Manning (2019).

### v3

@@ 3.2.1 Intrinsic-CFA
Intrinsic-CFA enhances the feature of query triplet by replacing its visual features of entities (i.e., subject and object). It consists of two steps: entity category selection and entity feature selection and replacement (c.f. Figure 4(a)). During training, we randomly select a query triplet from a batch of images, and randomly select one of its entity features as input. Then we put it into the Intrinsic-CFA module (c.f. Figure 3). Next, we detailedly introduce each step.
Entity Category Selection. This step is used for determining the category to which the entity feature of query triplet is replaced. Firstly, we propose a novel hierarchical clustering strategy to mine potentially fungible entity categories. Then we randomly select an entity category from the same cluster for the next entity feature selection. For the query triplet dog-laying on-bed in Figure 4, the category cat is selected from the same cluster as dog. The categories in the same cluster are common in the behavior patterns (e.g., they can be ridden), contexts (e.g., they can appear in a scene with a street) as well as semantics. Accordingly, three kinds of similarity are used to measure the common characters of two entity categories: pattern similarity, context similarity, and semantic similarity.
@@
(i) Pattern Similarity. It measures the overlap of the behavior patterns of two entity categories [39]. Figure 5(a) visualizes the common pattern between cat and dog (e.g., they all can be attached by the paw). Pattern similarity between two entity categories $c_{i}$ and $c_{j}$ is defined as:
$Si{m_{p}}({c_{i}},{c_{j}})$ $=\frac{{|{L_{s}}|}}{{{d_{out}}({c_{i}})+{d_{out}}({c_{j}})-|{L_{s}}|}}$ (6)
$+\frac{{|{L_{o}}|}}{{{d_{in}}({c_{i}})+{d_{in}}({c_{j}})-|{L_{o}}|}},$
where $|L_{s}|$ ($|L_{o}|$) is the number of common pred-obj (sub-pred) classes of the triplet whose sub (obj) class is $c_{i}$ or $c_{j}$. $d_{in}(c_{i})$ ($d_{out}(c_{i})$) is the number of incoming (outgoing) edges of the entity with class $c_{i}$ in the whole dataset.
(ii) Context Similarity. It measures the overlap of other entity categories in images of the two entity categories. Both cat and dog can appear in the same scene with entity categories, e.g., room in Figure 5(b). It is defined as:
$\small Si{m_{c}}({c_{i}},{c_{j}})=\frac{{|{L_{e}}|}}{{{d_{co}}({c_{i}})+{d_{co}}({c_{j}})-|{L_{e}}|}},\\$ (7)
where $|L_{e}|$ denotes the number of entity category intersections that appear in the same image with entity category $c_{i}$ or with $c_{j}$. $d_{co}(c_{i})$ is the number of entity instances in the whole dataset that appear in the same image as category $c_{i}$.
(iii) Semantic Similarity. It is the cosine similarity between two entity categories on semantic embedding [33] space:
[…]
The final similarity function $Sim({c_{i}},{c_{j}})$ used in clustering is the weighted sum of above three similarities. Complete clustering algorithm and results are in the appendix.
Entity Feature Selection and Replacement. This step selects an entity feature from the feature bank based on the selected entity category and replaces it with the query triplet. Specifically, we first update a new entity category of query triplet (e.g., cat-laying on-bed). Then, we select all triplet features with the same category as query triplet from feature bank as candidates (c.f. Figure 4(a)). And we utilize spatial restriction to filter unreasonable triplet features.
Spatial Restriction. In real augmentation, even if the two entity categories appear to be interchangeable, the positions of the subject and object after replacement may not make sense. For example, in VG dataset (c.f. appendix), though category hat and shoe are in the same cluster, directly replacing hat with shoe would place the shoe on the head and make it contrary to common sense.
To further ensure the rationality of the replacement of the two entity features, we use the cosine similarity of the relative spatial position of subject-object of the query triplet and the selected triplet from the feature bank as the restriction:
@@ 3.2.2 Extrinsic-CFA
Extrinsic-CFA aims to enrich the feature of the query triplet through the extrinsic information (i.e., the context formed by all the entities of an image) of context triplets. Different from the Intrinsic-CFA, we first select context triplet during training for computation efficiency. Extrinsic-CFA includes three steps: context triplet selection, query triplet feature selection, and feature augmentation (c.f. Figure 4(b)). Then, we will detailedly introduce each step of Extrinsic-CFA.
@@
Query Triplet Feature Selection. This step selects a query triplet with its features for the feature augmentation. Specifically, we select triplets with the same subject-object categories (e.g., pillow-bed) as the context triplet (e.g., pillow-on-bed) from feature bank as the candidate query triplets. Then, we implement the same spatial restriction as intrinsic-CFA to filter unreasonable triplets. Finally, we randomly select one of the reasonable triplets as the query triplet (e.g., pillow-laying on-bed), and its features are utilized for augmentation (c.f. Figure 4(b)).
Feature Augmentation. To reduce the impact on other triplets of the original image and enhance the extrinsic features of the tail predicate triplets, we perform mixup operation [47] between the selected query triplet (e.g., pillow-laying on-bed) and context triplet (e.g., pillow-on-bed), c.f. Figure 4(b). The mixup operation is written as:
$\tilde{\bm{v}}_{s}=\theta\bm{v}_{s}+(1-\theta){{\bm{v}}_{s}^{\prime}},$ (10)
$\tilde{\bm{v}}_{o}=\theta{\bm{v}}_{o}+(1-\theta){{\bm{v}}_{o}^{\prime}},$ (11)
$\tilde{\bm{u}}=\theta\bm{u}+(1-\theta){{\bm{u}}^{\prime}},$ (12)
where $\bm{v}_{s}^{\prime}$ ($\bm{v}_{s}$) and $\bm{v}_{o}^{\prime}$ ($\bm{v}_{o}$) denote the visual features of the subject and object of the query triplet (context triplet), respectively. The $\bm{u}^{\prime}$ ($\bm{u}$) denotes the visual feature of the uion box of the query triplet (context triplet). Similarly, predicate target ground-truth $r$ of the context triplet is mixed:
@@
After the mixup operation, the query triplet feature can be enhanced in context modeling (c.f. Figure 3) by leveraging the extrinsic features of the context triplet.
@@ 3.3 Training Objectives
Contrastive Loss. Due to the pattern changes between the entity features after mixup and original entity features, using only XE losses may result in performance drops in entity classification. To maintain the discriminative entity features after mixup, we further apply a contrastive loss [4]:
@@
[…]
Motifs+BPL-SA $1{}_{\textit{ICCV'21}}$ 29.7 31.7 50.7 52.5 41.2 16.5 17.5 30.1 31.0 23.8 13.5 15.6 23.0 26.9 19.8
[…]
Motifs+CFA (ours) 35.7 38.2 54.1 56.6 46.2 17.0 18.4 34.9 36.1 26.6 13.2 15.5 27.4 31.8 22.0
[…]
VCTree+BPL-SA $1{}_{\textit{ICCV'21}}$ 30.6 32.6 50.0 51.8 41.3 20.1 21.2 34.0 35.0 27.6 13.5 15.7 21.7 25.5 19.1
[…]
VCTree+CFA (ours) 34.5 37.2 54.7 57.5 46.0 19.1 20.8 42.4 43.5 31.5 13.1 15.5 27.1 31.2 21.7
[…]
Transformer+CFA (ours) 30.1 33.7 59.2 61.5 46.1 15.7 17.2 36.3 37.3 26.6 12.3 14.6 27.7 32.1 21.7
Table 1: Performance (%) of the SOTA trade-off SGG models on VG [20]. “Mean” is the average of mR@50/100 and R@50/100.
@@
[…]
Motifs+PPDL $_{\textit{CVPR'22}} 27$.9 32.2 33.3 15.8 17.5 18.2 9.2 11.4 13.5
[…]
Motifs+CFA (ours) 31.5 39.9 43.0 17.3 20.9 22.4 11.2 15.3 18.1
[…]
VCTree+CFA (ours) 31.6 39.2 42.5 21.5 26.3 28.3 10.8 15.1 17.9
[…]
Transformer+CFA (ours) 31.2 38.6 41.5 17.2 20.9 22.7 10.6 15.0 17.9
Table 2: Performance (%) of the SOTA tail-focused SGG models on VG [20]. $\ddagger$} means using the component prior knowledge.
@@
Table 3: Performance (%) of the SGG models on GQA [16].